A rxn of the form aA→products is second order with a rate constant of 0.225 M -1 s -1. If the initial conc of A is 0.293 M, what is the molar conc of A after 35.4 s?
1
Expert's answer
2017-10-13T15:14:07-0400
Integration of second-order rate law yields
1/[A] =1/[A_0 ] + kt;[A]=1/(kt+1/[A_0])
Time t, rate constant k and initial concentration [A_0] being known, one can calculate the final concentration of A :
[A]=1/(0.225 (M^(-1) s^(-1) )∙35.4(s)+ 1/0.293(M))=0.0879 (M)
Answer: 0.0879 (M)